OpenStudy (anonymous):

pAt=2+3t/5 Solve for t

OpenStudy (anonymous):

if it is 2 + (3t/5) if (2+3t)/5 pAt=2+3t/5 pAt= (2+3t)/5 pAt=10/5 +3t/5 5pAt= 2+3t pAt= (10+3t)/5 5pAt-3t=2 5pAt=10+3t t(5pA-3)=2 5pAt-3t=10 t=2/(5pA-3) t(5pA-3)=10 t=10/(5pA-3)